# Neil A. Chriss

### v2

@@ 1.1 Competition, implementation cost and equilibrium
What drives competition in trading and what are the traders competing for? We examine the situation in which two or more traders simultaneously believe that there is a catalyst that at some nearby future time will cause the price of a stock to rise to a new price called the target price [^1] This is a positive catalyst. There are obviously negative catalysts as well and the work here works identically for those. We will sometimes refer to these traders as the informed traders or just "the traders" when there is no risk of confusion. The difference between the current price and the target price represents a theoretical profit that is the motivation for traders to purchase the stock. Traders individually choose their target quantities to purchase and they all would like to capture as much of the theoretical profit as possible. Though these traders represent a minority of all traders who might trade in the stock, we model them as the only traders acting on the information related to the catalyst. Therefore their collective buying represents the marginal demand for the stock and puts upward pressure on its price. We presume all traders recognize this and therefore select the sizing and timing of their trades to minimize the price they part for the stock. Standing in the way are two forms of market impact, those effects that cause the price of the stock to move as a result of their collective trading.
@@
The first market impact effect is the persistent rise in price due to the collective and persistent demand for the stock from the informed traders. This persistent buying "leaks" into the market and causes other participants to re-rate the stock’s value upward. This is often referred to as permanent market impact or alpha-decay where in this context the former means permanent only until the catalyst occurs and the latter refers to the degradation of theoretical profit due to the trading. For a given trader, the cost of acquiring their target position is measured as the difference between the price per share of the stock immediately before trading begins and the average price paid is called the implementation cost of the trade.
@@
Each trader wishes to minimize their implementation cost by choosing the timing and sizing of their trades, the totality of is called a trading strategy. Each trader must select their strategy taking into account what their adversaries strategies are likely to be. One way to imagine this is that each trader is shifting portions of their target quantity backward and forward in time in order to balance two opposite considerations. Shifting trades earlier in time lowers later purchase prices by avoiding alpha-decay, but this comes at the expense of higher temporary impact. Determining the optimal mix depends on all the strategies but because traders are not coordinating wit h one another they need to select their strategies according to general principles. A Nash equilibrium (see [Nas50]) is the key concept in this regard.
@@ 1.2 The market strategy and equilibrium strategies
With this in hand a number of new results become available. First, total implementation cost of the market strategy, or equivalently, the sum of the cost to each trader is independent of how distributed and only depends on the number of traders and the temporary impact and alpha-decay parameters. This arises because there is a highly constrained relationship between the the individual strategies and the market strategy. Not only does this allow us to provide equilibrium strategies in closed-form (see Section 5) but also provide formulas for the implementation cost of each strategy, e.g., see Proposition 5.2. From this we immediately see another surprising fact: each traders implementation cost only depends on their target quantity as a fraction of all the other traders.
1.3 Centralized trading and disguising order flow
Modern firms such as multi-strategy hedge funds and mutual funds often have multiple traders who are trading in competition for the same stock despite working for the same firm. One can model this situation without regard to the firm itself, but we can also ask whether the firm should consider whether it can, in some, way consolidate trades to its advantage. In Section 6 we use the equilibrium strategies and cost formulas we derive to analyze the potential benefits of trade centralization from a game-theoretic standpoint, taking into account that there are still traders outside of the firm competing for the same stock. We find that if a firm opts to naively centralize and trade their order flow as a single, monolithic trade it may very well be worse off centralizing. Alternatively, a firm may strategically centralize its trades by disguising its true number of traders and in this case it has material opportunities to lower its implementation costs. In the former case, the reason firms will find it challenging to benefit from naive centralization is because in this case its orders will become large and vulnerable to front-running as described in Section 1. Naive centralization does provide benefits overall, but ironically they accrue to traders who are not in the firm. However, if a firm strategically centralizes, that is, if they take their consolidated order flow and work to mask the true number of traders and optimize, they there can be substantial improvements to implementation costs.
2 Detailed background information
@@
In this section we review the key concepts used through this paper in detail, repeating and enriching the overview in the introduction. For the most part these concepts were introduced in [Chr24a], however several, including the market strategy and trading coalitions, are new.
@@ 2.1 Catalysts, urgency and competition
This paper concerns trading in competition so the first thing to settle is why there is competition and how does this shape demand for a stock. Our idealized model begins with the concept of a catalyst that will occur at a fixed future date that has the power to significantly revalue the company and cause the stock price to significantly change. This is a fairly prevalent situation in equity markets world-wide; hedge funds and similar market participants do research to find catalysts and then build positions to capture the profit available from the price move. There are several important features of this setup that drive the need to explore a game-theoretic analysis.
@@ 2.2 Trading strategies
In order to studying trading strategies precisely we idealize the situation as follows. We assume there are $n$ traders who all start to trade at a fixed time $t=0$; they all commence trading at this time and measure their costs relative to the prevailing price at the time which is variously the reference price or arrival price. Prior to this time each trader each decides on their target quantity and we normalize all target quantities to 1. We write $\lambda_{i}$ for trader $i$’s target quantity as a fraction of the whole so that $\sum_{i=1}^{n}\lambda_{i}=1$. Thus target quantity represents the portion of the total quantity a trader seeks.
Each trader’s target quantity is regarded as a single order to buy the stock, and then the trader breaks that em parent order into subsidiary orders tied to it called child orders. A trading strategy is the precise scheduling of child orders so as to implement the parent order.
@@
  • A trading is called a position-building strategy if it starts with no stock (at $t=0$) and ends with a positive quantity of stock (at $t=1$);
  • We only consider position-building strategies and thus use the term trading strategy and position-building strategy analogously;
  • A position-building strategy whose target quantity is $1$ is called a unit strategy; and
  • When two or more traders are building a position in the same stock over the same interval we say they are building the positions in competition. This is the primary focus here and in [Chr24a].
  • When a trader is building a position in a stock, we will refer to this as taking up the trade toward acquiring the target quantity.
  • The position–building strategy is specified with two components: the target quantity $\lambda_{i}$ and the unit strategy $a_{i}$ where $i$ is the index of the trader.
@@ 2.3 A taxonomy of trading strategies
Since every trading strategy is represented by a function, it is possible to divide strategies into by the shape of their functions. Strategy shapes we discussed in [Chr24a] but for completeness we review the key points here. There are three categories of simple strategies, those whose second derivative is constant sign: risk-averse, risk-neutral and eager according to whether their second derivative is positive, zero or negative. Figure 1 gives a visual display of these:
@@
A different categorization of strategies is called aggressive strategies. These are strategies acquire more rapidly than the market-neutral strategies and do so in order to front-run other strategies in order to, in effect, their market impact their competitors (see Section 2.4 for more on market impact). Several examples of aggressive strategies are shown in Figure 2 and these are quite common in equilibrium settings.
@@
2.4 Implementation cost and market impact
@@
Here and throughout, we assume as usual that there is a small number of informed traders competing to buy a stock. We assume that each trader has selected a strategy and ask, what will it cost each trader to acquire their position? This cost has a specific meaning: it is the in excess of the reference price, that is, the price paid (per share) in excess of the prevailing price prior to the start of trading. At work is the assumption that the aggregate demand for liquidity caused by the traders outstrips supply, for example caused by their collective urgency, and the persistent supply-demand imbalance causes a gradual revaluation of the stock due to the presumed informativeness of the trades. In the whole, the price average execution cost of the acquisition is called the implementation cost and, of course, traders are expected to choose their strategies in such a way as to reduce their implementation cost.
We call the former component of market impact temporary impact and the latter as permanent impact or when we are referring to the erosion of theoretical profit, alpha-decay. Both temporary and permanent impact have been studied extensively from both theoretical points of view, including detailed analyses of order book dynamics and market microstructure. In the optimal execution and microstructure literature the persistent component.
Modeling market impact
We briefly explain how we model market impact in this paper. To start, at any given time trader $i$’s cumulative stock acquisition through time $t$ is given by its strategy, the function $\lambda_{i}a_{i}(t)$. Therefore their change in position between two nearby times $t_{1}$ and $t_{2}$ is $\lambda_{i}(a_{i}(t_{2})-a_{i}(t_{1}))$ which is the quantity of stock traded between those two times. Should we divide by $t_{2}-t_{1}$ we arrive at the rate of trading. We model the impact on the price of a stock due to immediacy, the temporary impact, as proportional to the rate of trading [^2] This is the standard linear temporary impact model studied in the optimal execution literature, see [AC97] and [AC01] for example..
As an example, if a trader sets units of time to one hour and trades at 5,000 units (e.g., dollars) per hour then we are are modeling the price impact as proportional to 5000 units per hour [^3] As we are using these models to select strategies the constant of proportionality is irrelevant. If we were using these models to accurately forecast price changes we would need to precisely calibrate the model for that purpose. Since the trader also is buying the 5,000 units, the total increased price paid is proportional to $000\times 5,$ 000 × 5 , 000 units. Thus the impact on implementation cost is quadratic. Finally, it is important to note that we will be modeling the aggregate impact that all of the traders trading in competition have on the market, and any one in particular.
We model permanent impact quite differently based on the cumulative quantity acquired. In this case, our model reflects the notion that as the informed traders collectively trade, the total quantity they acquire is proportional to the level of information the collectively "leak" into the market and the impact on price is proportional to that. We model this as permanent impact is proportional to the cumulative stock purchased by all of the traders through a given time. We write $\kappa$ for the constant of proportionality and call it the market impact parameter; $\kappa$ controls the relative contribution of permanent impact versus temporary impact.
The market strategy and market impact
We write $m(t)=\sum_{i=1}^{n}\lambda_{i}a_{i}(t)$ for the implied trading strategy due to the aggregate trading of all the competitive traders. We call this the market strategy (this is neither the market nor a strategy however it is elaborated on extensively later in this paper). Because the competitive traders represent the marginal supply and demand for the stock their collective trading is the key driver of market impact. We model the instantaneous cost of trading as proportional to the market strategy’s rate of trading and the permanent impact as proportional to the cumulative trading by the market strategy. Both of these model the change in stock price, the former relative to the prevailing market price and the latter relative to the reference price. Adding these up, says that the total cost at time $t$ is proportional to:
$C_{i,\lambda_{i}}(t)\equiv\dot{m}(t)\,\lambda_{i}\dot{a}_{i}(t)+\kappa\,\lambda_{i}\,m(t)\dot{a}_{i}(t)$ (2.1)
@@
This value represents the cost relative to every trade occurring at the prevailing price without any permanent impact due to trading and only up to a constant multiple. Thus it is not a measure of the actual cost of trading at time $t$; however, since our aim is to find optimal, that is lowest cost, strategies, we are only interested in cost up to a constant.
@@ 2.5 Minimizing implementation cost
The instantaneous cost function in Eq. 2.1, $C_{i,\lambda_{i}}$ takes into account the trading of all traders and both the temporary impact and alpha-decay cost of trading. The implementation cost to trader $i$ is the integral from $t=0$ to $t=1$ of the cost function:
$\text{Cost}(a_{\lambda_{i}},m)=\int_{0}^{1}C_{i,\lambda_{i}}(t)\,\mathrm{dt}=\int_{0}^{1}\left(\dot{m}(t)+\kappa\,m(t)\right)\lambda_{i}a_{i}(t)\,\mathrm{dt}$ (2.2)
Equation (2.2) is the total implementation cost to trader $i$ of trading in competition with traders $i=1,\dots,n$ whose market strategy is $m=\sum\lambda_{i}a_{i}$. This is the juncture at which game theory enters the scene because each trader would like to minimize their implementation cost with respect to the others. Thus, each trader needs to make a choice based on what the others are likely to do.
@@
A tool we will use repeatedly to find strategy that minimize implementation cost is the Euler-Lagrange equation. This states that a function $a_{i}:[0,1]\to\mathbf{R}$ that minimizes the above integral satisfies the following linear partial differential equation:
$\frac{\partial}{\partial t}\frac{\partial C_{i,\lambda_{i}}}{\partial\dot{a}_{i}}-\frac{\partial C_{i,\lambda_{i}}}{\partial a_{i}}=0$ (2.3)
Since our aim is to find Nash equilibria, our aim is therefore to find strategies $\lambda_{1}c_{1},\dots,\lambda_{n}c_{n}$ such that each trader’s strategy simultaneously minimizes their cost function, which clearly involves finding the simultaneous solution to Eq. 2.3 for all $i$.

### v3

@@ 1.1 Competition, implementation cost and equilibrium
What drives competition in trading and what are the traders competing for? We examine the situation in which two or more traders simultaneously believe that there is a catalyst that at some nearby future time will cause the price of a stock to rise to a new price called the target price [^1] This is a positive catalyst. There are obviously negative catalysts as well and the work here works identically for those. We will sometimes refer to these traders as the informed traders or just "the traders" when there is no risk of confusion. The difference between the current price and the target price represents a theoretical profit that is the motivation for traders to purchase the stock. Traders individually choose the quantity they wish to trade, called their target quantity.Each would like to capture as much of the theoretical profit as possible.
Though these traders represent a minority of all traders who might trade in the stock, we model them as the only traders acting on the information related to the catalyst. Therefore their collective buying represents the marginal demand for the stock and puts upward pressure on its price. We presume all traders recognize this and therefore select the sizing and timing of their trades to minimize the price they part for the stock. Standing in the way are two forms of market impact, those effects that cause the price of the stock to move as a result of their collective trading.
@@
The first of the two market impact effects is the persistent rise in price caused bu the collective and persistent demand for the stock from the informed traders. This persistent buying "leaks" into the market and causes other participants to re-rate the stock’s value upward. This is often referred to as permanent market impact or alpha-decay where in this context the former means permanent only until the catalyst occurs and the latter refers to the degradation of theoretical profit due to the trading. For a given trader, the cost of acquiring their target position is measured as the difference between the price per share of the stock immediately before trading begins and the average price paid is called the implementation cost of the trade.
@@
Each trader wishes to minimize their implementation cost by choosing the timing and sizing of their trades, the totality of which is called their trading strategy. Traders select their strategies taking into account what their adversaries’ strategies are likely to be. One way to think about what "choosing a strategy" means is to imagine that each trader shifts a portion of their target quantity backward and forward in time in order to balance two opposite considerations. Shifting trades earlier in time lowers later purchase prices by getting ahead of alpha-decay, but to do so comes at the cost of higher temporary impact. Determining the optimal mix depends on all the strategies but because traders are not coordinating wit h one another they need to select their strategies according to general principles. A Nash equilibrium (see [Nas50]) is the key concept in this regard.
@@ 1.2 The market strategy and equilibrium strategies
With this in hand a number of new results become available. First, the total implementation cost of the market strategy, or equivalently, the sum of the cost to each trader, is independent of what share of the total target quantity each trader seeks is distributed. It only depends on the number of traders, the temporary impact and alpha-decay parameters. This arises because there is a highly constrained relationship between the the individual strategies and the market strategy. Not only does this allow us to provide equilibrium strategies in closed-form (see Section 5) but also provide formulas for the implementation cost of each strategy, e.g., see Proposition 5.2. From this we immediately see another surprising fact: each traders implementation cost only depends on their target quantity as a fraction of all the other traders.
1.3 Centralized trading and masking a firm’s number of traders
Modern investment firms such as multi-strategy hedge funds often have multiple traders who are trading in competition for the same stock despite working for the same firm. We can model this situation without regard to the firm itself, that is, we can treat each trader as independent and detached from the others. We can also ask whether the firm can consolidate its trades to its advantage, that is to effect lower implementation costs.
In Section 6 we use the equilibrium strategies and cost formulas we derive in previous sections to analyze the potential benefits of trade centralization. The key observation with respect to centralization is that prior discussions of trading in competition assume that each trader acts completely independently with no communication or coordination with the others. However, when a firm employs a subset of the traders in competition, it is possible for the firm to modify some aspects of the game traders are engaged in, in particular the number of traders in competition for the given stock. The firm can first centralizing all its trade and then mask the true number of traders it has an "represent" to the market that it has any number it likes by simply consolidating or splitting up its orders, never changing its order quantity.
We study this in detail in Section 6 using the various results regarding equilibrium trading costs. These allow us to compute the total cost to all traders in competition and it is a curious fact (though a familiar one to game-theorists) that as the total number of traders grows, the costs of trader in equilibrium grow as well, see for example Figure 13. Because some number of traders work for the same firm, the firm has an opportunity to "change the game" through centralization, specifically by changing the total number of traders in competition. This itself might lead one to think that a firm should therefore simply centralize their trades and trade them as a single trader, which indeed would minimize the total number of traders "in the game". However, there is another consideration, this time that for a given trader who is trading a relative large share of the total quantity, the fewer players there are, the greater a share of total costs that trader will incur (see Figure 5 and the discussion there).
Therefore the must be an optimal balance with regard to centralization. The optimal number of traders to "represent" is surprisingly simple. The optimal number of traders to represent is approximately the number of traders who do not work for the firm. This is outlined in Section 6.4 and the result itself is Eq. 6.14. Perhaps the most surprising result of all is that the optimum number of traders itself does not depend on any of the other quantities involved, such as target quantities or alpha-decay.
2 Background information and key definitions
@@
In this section we review the key concepts used through this paper in detail, repeating and enriching the overview in the introduction. For the most part these concepts were introduced in [Chr24a], however several, including the market strategy, are new.
@@ 2.1 Catalysts, urgency and competition
This paper concerns trading in competition so the first thing to settle is why there is competition and how does this shape demand for a stock. Our idealized model begins with the concept of a catalyst that will occur at a fixed future date that may cause a significant revaluation of the stock, that is, its price may significantly rise immediately after the catalyst. This is a fairly common situation in equity markets: hedge funds and similar market participants conduct research to find catalysts and then build positions to capture the theoretical profit available from the potential price moves. There are several important features of this setup that drive its exploration from a game-theoretic point of view.
@@ 2.2 Trading strategies
We study an idealized situation and assume there are $n$ traders seeking to buy the same stock starting simultaneously at a time designated as time $t=0$. The traders are referenced by the indexes $i=1,\dots,n$. All measure their cost of acquiring the stock relative to the prevailing price at $t=0$. This price is called the reference price or the arrival price. At some time prior to $t=0$ each trader decides how much stock to acquire, their target quantity. We write $\lambda_{i}$ for trader $i$’s target quantity and normalize their total so that $\sum_{i=1}^{n}\lambda_{i}=1$, and refer to the $\lambda_{i}$’s as fractional target quantities.
At the start of trading we regard each trader’s intention to buy the stock as a single order to buy, called the parent order. This will be broken up into a number of smaller orders, the child orders, over time. The details of how a parent order is broken up into child orders is called the trading strategy. We formalize these ideas as follows:
@@
  • A trading strategy is called a position-building strategy if it starts with no stock at $t=0$ and ends with a positive quantity of stock (at $t=1$). We will generally discuss position-building strategies in this paper and therefore we will use the term trading strategy to mean position-building strategy unless otherwise stated;
  • A position-building strategy whose $a_{i}(t)$ such that $a_{i}(1)=1$ is called a unit strategy;
  • We will assume each trader (say $i$) has a fixed quantity of the stock they wish to acquired called the target quantity. After we normalize the sum of all target quantities to one, we will denote trader $i$’s target quantity by $\lambda_{i}$ and refer to it as the fractional target quantity;
  • When two or more traders are building a position in the same stock over the same interval we say they are building the positions in competition. This is the primary focus here and in [Chr24a];
  • The activity of multiple traders competing for the same stock will be referred to as the trade, and each is said to be in the trade or taking up the trade; and
  • Each trading strategy is specified by two components: a fractional target quantity $\lambda_{i}$ and a unit strategy $a_{i}$ where $i$ is the index of the trader.
@@ 2.3 A taxonomy of trading strategies
Since every trading strategy is represented by a function of time, it is possible to categorize strategies by the shape of their graphs. Strategy shapes were discussed in [Chr24a] but for completeness we review the key points here. There are three categories of simple strategies, that is, strategies whose second derivatives have constant sign: risk-averse, risk-neutral and eager according to whether their second derivative is positive, zero or negative. Figure 1 gives a visual display of these:
@@
Eager strategies acquire more stock more rapidly than the market-neutral strategies and do so in order to front-run other strategies, meaning, in effect, to avoid the permanent market impact of their competitors (see Section 2.4. Some eager strategies will overbuy and at an interim time begin to sell. We give examples of eager strategies that overbuy, which we will call aggressive, in Figure2.
@@
2.4 Market impact and implementation cost
@@
Here and throughout, we assume as usual that there is a small number of informed traders competing to buy a stock. We assume that each trader has selected a strategy and ask, what will it cost each trader to acquire their position? This cost has a specific meaning here. It is the price in excess of the reference price, that is, the price paid (per share) in excess of the prevailing price prior to the start of trading. At work are two forces, one relating to the transient impact of liquidity demand and the other related to the collective information content of the trades. The first is based on the assumption that at any given moment the aggregate demand for liquidity in the stock caused by the traders may outstrip available supply causing traders to pay a premium to have their trades executed. The second is based on the idea that when traders trade toward a catalyst the market is able to spot to some degree this activity and extract some information value from it which causes the price to persistently rise as a consequence of the trading. We work out the details of this now.
Trading strategies, derivatives and the market strategy
We use the same temporary and permanent impact models as in [AC97], [AC01] and [Chr24a]. In brief, temporary market impact is the immediate and incremental cost of trading due to transient, liquidity driven effects. This is modeled as having a per unit impact on price proportional to the net rate of trading, where net refers to the total over all traders trading in competition. Permanent impact, on the other hand, is an effect due to the cumulative level of trading over the life of a trade and this is modeled as having a per unit impact on price proportional to the cumulative quantity of stock acquired throughout the life of a trade. To explore these further we need to define precisely what we mean by rate of trading.
To begin, we lay out our standard notation and setup. There are $n$ traders in competition identified by indexes $i=1,\dots,n$. We write $\lambda_{i}$ for the fractional target quantity of trader $i$. This is the fraction of the total target quantity over all traders that trader $i$ will trade so that $\sum_{i=1}^{n}\lambda_{i}=1$ and therefore total quantity is normalized to one. We write $a_{i}(t)$ for the unit part of $i$’s trading strategy, and therefore $a_{i}(t)=0,a_{i}(1)=1$ and $\lambda_{i}\,a_{i}(t)$ is formally the trader’s strategy. For a given strategy we write $\dot{a}_{i}(t)$ for the derivative of the strategy with respect to time. This represents the rate of trading in the stock. It is instantaneous rate of trading and may be regarded as an instantaneous measure of how much trader $i$ is trading at a given time.
We also write
$m(t)=\sum_{i=1}^{n}\lambda_{i}a_{i}(t)$ (2.1)
for the market strategy. We use this anthropomorphism in the same spirit as when one says "the market rallied", despite there being no collective volition involved. We note that
$\dot{m}(t)=\sum_{i=1}^{n}\lambda_{i}\dot{a}_{i}(t)$ (2.2)
is the net rate of trading of all traders in competition. With these concepts in hand we can formally define temporary and permanent market impact.
Temporary and permanent impact
As usual we are concerned with $n$ traders in competition for a specific stock and we focus our attention on the trading that occurs between time $t=0$ (the start time) and $t=1$ (the completion time). Note that both time and trading quantity are normalized to one unit, which means in effect that we have erased all of the details concerning the level of trading relative to the level of available liquidity. We will take this point up later in the paper.
As already noted we are interested in each trader’s implementation cost and this arises from two effects acting in tandem on the execution price of the stock and we reiterate the fact that as in [Chr24a] both temporary and permanent impact arise from the collective trading of all the traders in competition for the stock. In this way of looking at things, each time a trader executes a trade there are two things impacting the incremental price paid due to temporary impact:
  • The trader’s rate of trading: for, say, trader $i$, the rate of trading at time $t$ is given by $\lambda_{i}\dot{a}_{i}(t)$. This is effectively the quantity of traded over an infinitesimal duration of time. One way to see this is to note that integral of $\lambda_{i}\,\dot{a}_{i}(t)$ over the interval $[t,t+\Delta]$ is precisely the change in quantity acquired over that interval.
    $\int_{t}^{t+\Delta t}\lambda_{i}\,\dot{a}_{i}(t)\,\mathrm{dt}=\lambda_{i}\,\left(a(t+\Delta t)-a(t)\right)$
    Thus $\lambda_{i}\,\dot{a}_{i}(t)$ acts as a stand-in for the quantity traded; and
  • The aggregate (net) rate of trading for all traders who have taken up the trade: we model the incremental supply or demand for the stock as the trading pushing the price. We already defined the market strategy $m(t)=\sum_{i=1}^{n}\lambda_{i}a_{i}(t)$ and we therefore use $\dot{m}(t)$ to model the net quantity of traded by all traders over a short period of time. We model the impact on the price of the stock at time $t$ as proportional to $\dot{m}(t)$.
The product of $\lambda_{i}\dot{a}_{i}(t)$ and $\dot{m}(t)$ is therefore the product of the infinitesimal quantity traded and the impact on price per share and this is how we model the total cost of trading due to temporary impact at a given time $t$.
[…]
We model the cost of trading to trader $i$ due to persistent demand for the stock in a similar way:
[…]
In Eq. 2.4 we are expressing that the cost of trading due to permanent impact is proportional to the product of the rate of trading and the cumulative quantity acquired by all traders. This in effect says that there is a relationship between the quantity these traders acquire and how much information leaks into the market. Since our concern is always to minimize the total cost of trading, we focus on the sum of the temporary and permanent components of market impact. We summarize this in the function $C_{i}(t)$ for trader $i$ defined with $\lambda_{i}$ implicit as follows:
$C_{i}(t)=\dot{m}(t)\,\lambda_{i}\dot{a}_{i}(t)+\kappa\,\,m(t)\lambda_{i}\dot{a}_{i}(t)$ (2.5)
@@
Note that we are only interested in the value of $C_{i}$ up to a multiplicate constant in that all of our uses for it are in cost minimization problems. In particular this means that this is not a measure of the actual cost of trading at time $t$; however, since our aim is to find cost optimal strategies, that is, the lowest cost, we are only interested in cost up to a constant. In particular this allows us to always set the constant of proportionality in temporary impact to one.
The function $C_{i}$ is called the instantaneous cost function for trader $i$ and it is the object of interest for finding equilibria in trading across multiple traders. Each trader would like to minimize their cost, but since each cost function involves $m$ this meet the criteria mentioned above of requiring each trader to take into account what all other traders are likely to do.
@@ 2.5 Minimizing implementation cost
The key assumption in this paper is that traders will choose their strategies to minimize their implementation costs and have to do so based on some rational notion of what other traders will do. This will lead to the notion of Nash equilibrium stated below but first we examine implementation cost in more detail. The instantaneous cost function in Eq. 2.5, $C_{i}$ takes into account the trading of all traders as the sum of temporary impact and permanent impact due to alpha-decay. The implementation cost to trader $i$ is therefore the integral from $t=0$ to $t=1$ of the cost function:
$\text{Cost}(\lambda_{i}a_{i},m)=\int_{0}^{1}C_{i}(t)\,\mathrm{dt}=\int_{0}^{1}\big{(}\dot{m}(t)+\kappa\,m(t)\big{)}\lambda_{i}a_{i}(t)\,\mathrm{dt}$ (2.6)
@@
Of note, (2.6) is the total implementation cost to trader $i$ of trading in competition with traders $i=1,\dots,n$ whose market strategy is $m=\sum\lambda_{i}a_{i}$ up to a multiplicative constant. This constant is irrelevant, however, as our purpose is to minimize costs and the primary tool we use is the calculus of variations. Specifically the tool we use to find cost-minimizing strategies is the Euler-Lagrange equation. This states that a function $a_{i}:[0,1]\to\mathbf{R}$ that minimizes Eq. 2.6 satisfies the following linear partial differential equation, called the Euler-Lagrange equation:
$\frac{\partial}{\partial t}\frac{\partial C_{i}}{\partial\dot{a}_{i}}-\frac{\partial C_{i}}{\partial a_{i}}=0$ (2.7)
Note that if $C_{i}$ satisfies this equations then so does $\zeta C_{i}$ for any constant $\zeta\neq 0$ and it is in this sense that we only need to know the cost functions up to a multiplicative constant.
We now introduce the definition of Nash equilibrium (see [Nas50]) which is the fundamental object of study in this paper.
Definition 2.1 (Nash equilibrium).
Consider the situation of $n$ traders in competition for the same stock each with fractional target quantities $\lambda_{1},\dots,\lambda_{n}$. Assume that in this context the market impact parameter is $\kappa$ then an equilibrium is a set of trading strategies $a_{1},\dots a_{n}$ and corresponding market strategy $m=\sum_{i=1}^{n}\lambda_{i}a_{i}$ with cost corresponding cost functions $C_{i}(t)=\lambda_{i}\dot{m}(t)\,a_{i}(t)+\lambda_{i}\kappa\,m(t)\,\dot{a}_{i}(t)$ such that each of
$\text{Cost}(\lambda_{i}a_{i},m)=\int_{0}^{1}C_{i}(t)\,\mathrm{dt}$ (2.8)
is simultaneously minimal. $\blacksquare$
Section 5 gives closed-form formulas for equilibrium strategies in general which generalizes the corresponding formulae for two-trader and $n$-trader symmetric equilibria in [Chr24a].
2.6 A note on time and quantity normalization
Trader $i$ with trading strategy $a_{i}(t)$ and target quantity $\lambda_{i}$ will incur a total implementation cost equal to the integral of $C_{i}(t)$:
$\int_{0}^{1}\dot{m}(t)\lambda_{i}\dot{a}_{i}(t)+\kappa m(t)\lambda_{i}\dot{a}_{i}(t)\,\mathrm{dt}$ (2.9)
and the Euler-Lagrange equation Eq. 2.7 tells us that a strategy $a_{i}(t)$ is cost-minimizing for any non-zero multiple $\zeta$ of $C_{i}$. In particular we can divide by $\lambda_{i}$ and write
$\int_{0}^{1}\dot{m}(t)\,\dot{a}_{i}(t)+\kappa\,m(t)\dot{a}_{i}(t)\,\mathrm{dt}$ (2.10)
and find the path $a_{i}(t)$ that minimizes Eq. 2.10 and this will yield the same strategy cost-minimizing strategy $a_{i}(t)$. Given this we make several points to try explain what the normalizations mean in this context.
  • Because Eq. 2.10 has the same cost-minimizing strategy as Eq. 2.9, $\dot{m}(t)\,\dot{a}_{i}(t)+\kappa\,m(t)a_{i}(t)$ is an equally valid cost function. The target quantity of each trader is still "present" but it is buried in $m(t)=\sum_{i=1}^{n}\lambda_{i}a_{i}(t)$;
  • At times we will continue to use the cost function $C_{i}(t)$ as opposed to the integrand in Eq. 2.10 (leaving in the $\lambda_{i}$ terms) because some of the machinations are more intuitive in this case;
  • We could easily have tacked a parameter $\iota$ onto temporary impact and thus made the cost function
    $\iota\,\dot{m}(t)\dot{a}_{i}(t)+\kappa\,m(t)\dot{a}_{i}(t)$ (2.11)
    and the Euler-Lagrange equation will continue to yield the same cost-minimizing strategy. While Eq. 2.11 offers a more precise description of both the temporary and permanent impact components by highlighting that each has its own degree of responsiveness to rate of trading and total trading respectively, it is mathematically more awkward so we implicitly multiple through by $\iota^{\sm 1}$ and in practice maintain a factor of 1 in the $\dot{m}(t)\dot{a}_{i}(t)$ term.
The impact of time and quantity normalizations
As all integrals are from 0 to 1 and all target quantities sum to 1, it is worth asking what may be lost in making these normalizations. From a purely mathematical standpoint, the answer is nothing. However, in terms of the interpretation of the terms, something is indeed lost.
To understand this, consider the following example. Suppose there are two similar trading situations both with three traders respectively trading 50%, 30% and 20% of the total target quantity. However, consider the additional data that the two cases differ only in their time frames:
  • Case 1: the actual time frame is 20 days (four weeks) and the actual target quantity is 10,000,000 shares; and
  • Case 2: the actual time frame is 2 days and (again) the actual target quantity is 10,000,000 shares.
In the models in this paper and in [Chr24a], case 1 and case 2 appear to be identical. Both would normalize the span of time over which the stock is purchase to the mathematically convenient interval $[0,1]$. Yet clearly these two cases are very different. To begin with, compressing the trading of 10,000,000 shares of stock from 20 to 2 days requires ten times the volume of stock to be traded on average at any given time. Given the temporary impact term sets the price impact of trading to be proportional to the instantaneous rate of trading at any given time, then on an all else equal basis, the temporary impact of trading a given strategy would be ten times greater in case 2 (two days) versus case 1 (twenty days). The question, then, is to what extent to the parameters of market impact change when time frames are compressed or dilated?
While this is not a question for this paper, it is worth asking whether in case 2 with ten times as much trading flowing through the market, whether the temporary impact parameter would change. Perhaps it would be be 10 times greater, which when re-normalized, the implementation cost would be
$\dot{m}(t)\dot{a}_{i}(t)+\frac{1}{10}\kappa\,m(t)\dot{a}_{i}(t)$ (2.12)
However, we do not know if this is correct because we also have to ask whether the level of alpha-decay expands or contracts when trading occurs over shorter durations. There are two potential considerations here. First, over a shorter period of time there is less opportunity for the market to become aware of informativeness of the trades. On the other hand, given the volume from the traders in competition is on average ten times greater, it may it stands to reason that there is more information being conveyed on average per unit of time. How do these two forces reconcile? While we will not pursue these questions here, it is necessary to point them out to properly understand their relation to real-world trading situations. Finally, we remark that while the above two cases involved "same quantity, different time frame", a similar set of questions clearly arise when dealing with same time frame, different quantity scenarios.